OpenStudy (anonymous):

If 0.11k+1.5=2.49,what does k=?(please show steps)

OpenStudy (anonymous):

9

OpenStudy (anonymous):

.11k + 1.5 = 2.49 -1.5 -1.5 ----------------- .11k = .99 DIVIDE BY .11 k = .9

OpenStudy (anonymous):

0.11k = 2.49-1.5 0.11k = 0.99 k=0.99/0.11 k=9

OpenStudy (anonymous):

i meant 9. sorry